# SceneSplat++: A Large Dataset and Comprehensive Benchmark for Language Gaussian Splatting

### v3

@@ Abstract
3D Gaussian Splatting (3DGS) serves as a highly performant and efficient encoding of scene geometry, appearance, and semantics. Moreover, grounding language in 3D scenes has proven to be an effective strategy for 3D scene understanding. Current Language Gaussian Splatting line of work falls into three main groups: (i) per-scene optimization-based, (ii) per-scene optimization-free, and (iii) generalizable approach. However, most of them are evaluated only on rendered 2D views of a handful of scenes and viewpoints close to the training views, limiting their ability and insight into holistic 3D understanding. To address this gap, we propose the first large-scale benchmark that systematically assesses these three groups of methods directly in 3D space, evaluating on 1060 scenes across three indoor datasets and one outdoor dataset. SceneSplat-Bench results demonstrate a clear advantage of the generalizable paradigm, particularly in relaxing the scene-specific limitation, enabling fast feed-forward inference on novel scenes, and achieving superior segmentation performance. We further introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising around 49K diverse indoor and outdoor scenes obtained from multiple sources, with which we demonstrate the generalizable approach could harness strong data priors. Our codes, benchmark, and datasets are available at SceneSplat-Bench and SceneSplat-49K.
@@ 1 Introduction
The field of 3D computer vision is highly focused on capturing the geometry WITKIN198117; Kazhdan2006; Hirschmuller2008StereoDepth; Newcombe2011KinectFusion; schoenberger2016sfm; Park2019DeepSDF; shi4568335city; ren2024bringing and the visual appearance 10.5555/539405; 791235; 5226635; mildenhall2020nerf; kerbl3Dgaussians of a scene, as well as understanding its content 8099499; wu2024ptv3; peng2023openscene; zhu2023pointclip; sarkar2025crossover. Recently, 3D Gaussian Splatting (3DGS) kerbl3Dgaussians has emerged as the most desirable 3D representation due to its unique ability to jointly encode geometry, appearance, and scene understanding properties Ye2024GSgroupping; Shi2024LEG; jiang2024open in a compact form that can be effectively optimized from 2D posed images ye2025gsplat. Furthermore, vision-language reasoning represents the most promising direction for 3D scene understanding rozenberszki2022language; ding2022pla; yang2024regionplc; ma2025cityloc; jiang2024open; wu2023mars; halacheva2025gaussianvlm, as it bridges the visual and geometric attributes of a scene with the language we use to define, describe, and reason about concepts. Therefore, this work focuses exclusively on vision-language scene understanding with 3DGS.
@@
The most relevant methods for language Gaussian Splatting (LGS) can be categorized into three groups. The first two groups begin by extracting 2D features from all training images using vision-language foundation models (e.g., CLIP Radford2021CLIP). The first group then performs gradient-based per-scene optimization Qin2024LangSplat; Zuo2024FMGS; zhou2024feature3DGS; Wu2024OpenGaussians; QU2024GOI, assigning feature vectors to each 3D Gaussian primitive and optimizing them such that their renderings align with the corresponding 2D feature maps. The second group also operates on a per-scene basis but adopts an optimization-free approach joseph2024GWfeatureBP; marrie2024ludvig; cheng2024occam. Instead of iterative refinement, these methods directly lift 2D features to 3D primitives via weighted feature aggregation schemes. Lastly, we consider the generalizable approach. To date, only a single method has been trained on 2K indoor scenes li2025SceneSplat7k, learning to directly predict a vision-language feature vector for each 3D Gaussian primitive. Although 2D vision-language features are required during training, inference relies solely on 3D Gaussians, eliminating the need to prepare additional 2D feature maps. Furthermore, all features are produced in a single forward pass.
@@
While these methods represent important advances in integrating vision-language reasoning with 3DGS, their evaluation protocols still suffer from several critical limitations. First, most methods are evaluated on only a small number of selected scenes Qin2024LangSplat; Zuo2024FMGS; zhou2024feature3DGS; Wu2024OpenGaussians; QU2024GOI; joseph2024GWfeatureBP; cheng2024occam. As a result, such evaluations carry a high risk of producing scene-specific rather than generalizable conclusions. Moreover, the absence of a standardized benchmark further limits systematic and comparable evaluation. Second, most methods are evaluated close to the training views Qin2024LangSplat; marrie2024ludvig; Wu2024OpenGaussians. Consequently, there is no guarantee that the results generalize well to novel viewpoints. Finally, most evaluations are conducted in 2D rather than directly in 3D space Qin2024LangSplat; Zuo2024FMGS; zhou2024feature3DGS; Wu2024OpenGaussians; QU2024GOI; joseph2024GWfeatureBP; marrie2024ludvig; cheng2024occam. However, 3D scene understanding fundamentally concerns performance in 3D space, which cannot be fully inferred from 2D projections. Therefore, we introduce SceneSplat-Bench, an evaluation benchmark for 3DGS vision-language scene understanding that includes 1060 scenes spanning 325 unique semantic classes. It evaluates performance in 3D for each segment of the scene, and we use it to assess representative methods from all three groups, as presented in Tab.˜1.
@@
Among the evaluated approaches, the generalizable 3D scene understanding paradigm demonstrates the strongest potential. It eliminates the need for substantial per-scene computation at inference and enables a single forward pass per 3D scene. In addition, this approach enables the application of 3D foundation models for scene understanding without the need for task- or scene-specific retraining. This concept aligns with established practices in 2D computer vision, where pretrained foundation models have been widely adopted He_2016_CVPR; dosovitskiy2020vit; oquab2023dinov2; ravi2024sam2; Radford2021CLIP; liu2023llava. Foundation models tailored to 3DGS have already emerged across key areas, including reconstruction charatan23pixelsplat; chen2024mvsplat; gslrm2024; ziwen2024llrm, scene understanding li2025SceneSplat7k, and generation xiang2024structured; roessle2024l3dg. Lastly, benchmark results demonstrate the state-of-the-art performance of the generalizable approach, highlighting its capacity to leverage data priors and extract meaningful vision-language features—effectively mitigating the noise inherent in weak supervision.
@@
To facilitate the development of generalizable 3DGS scene understanding, we introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising diverse indoor and outdoor scenes collected from multiple sources. Existing large-scale 3DGS datasets primarily focus on single objects ma2024shapesplat; liu24uco3d; wei20253daffordsplatefficientaffordancereasoning, whereas scene-level datasets li2025SceneSplat7k remain limited in scale. To the best of our knowledge, SceneSplat-49K represents the most extensive open-source dataset of complex, high-quality 3DGS reconstructions at the full-scene level. The preparation of the dataset required approximately 891 GPU days and considerable human involvement. Furthermore, we demonstrate that training a generalizable 3DGS scene understanding method on a larger subset from SceneSplat-49K leads to improved performance, achieving state-of-the-art results on the benchmark.
@@
Methods
Generalizable
Optimization-
free
Runtime
/ Scene
Benchmark
Dataset
Scene
Number
Class
Number
@@ Methods
LangSplat Qin2024LangSplat 621 min LERF kerr2023lerf, 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
FMGS Zuo2024FMGS 76 min LERF kerr2023lerf, 3DOVS liu2024weaklysupervised3dopenvocabulary 11 123
Feature3DGS zhou2024feature3DGS 235 min Replica straub2019replica 4 7
OpenGaussian Wu2024OpenGaussians 297 min LERF kerr2023lerf, ScanNet dai2017scannet(10scenes) 14 76
GOI QU2024GOI 252min Mip-NeRF360 barron2021mipnerf, Replica straub2019replica 8 30
@@
Gradient-Weighted
3DGS
 joseph2024GWfeatureBP
4 min LERF kerr2023lerf, 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
LUDVIG marrie2024ludvig 5.6 min LERF kerr2023lerf, SPIn-NeRF spinnerf, NVOS ren2022neural 21 80
OccamLGS cheng2024occam 3.2 min LERF kerr2023lerf, 3DOVS liu2024weaklysupervised3dopenvocabulary 9 88
@@
SceneSplat li2025SceneSplat7k 0.24 min ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++,Matterport3D chang2017matterport3d 732 321
SceneSplat-Bench
ScanNet dai2017scannet, ScanNet++ yeshwanth2023scannet++,
Matterport3D chang2017matterport3d, HoliCity zhou2021holicity,
1060 325
@@ 2 Related Works
Gaussian Splatting Datasets. Most existing 3DGS datasets remain object-centric, synthetic, or limited in scale and scope. Moreover, recent works charatan23pixelsplat; roessle2024l3dg; ziwen2024llrm; chen2024splatformer showcase 3DGS generalization, yet do not fully release their datasets, thus limiting reproducibility and scalability. Among publicly available datasets, ShapeSplat ma2024shapesplat transforms 65K object-level CAD models into 3DGS. 3DAffordSplat wei20253daffordsplatefficientaffordancereasoning focuses on functionality, providing 23K object instances labeled with affordances across 21 categories. uCO3D liu24uco3d expands the object-centric landscape with over 1,000 annotated categories and 360 coverage. Furthermore, scene-level 3DGS remains underexplored, with SceneSplat li2025SceneSplat7k being the only available dataset providing 6.8K indoor scenes to support open-vocabulary 3D understanding, though its scope is limited to the indoor domain. In summary, existing datasets fall short of enabling generalizable, multi-modal learning across diverse scenes. To address this, we introduce SceneSplat-49K, a large-scale dataset of 49K 3DGS scenes spanning indoor and outdoor environments, enriched with vision-language embeddings to support scene-level understanding.
@@
Open-Vocabulary Scene Understanding. Recent advancements in open-vocabulary models have significantly expanded the capabilities of vision-language understanding. Foundation models zhang2022dino; oquab2023dinov2; kirillov2023segment; ravi2024sam; radford2021learning; zhai2023sigmoid; tschannen2025siglip enabled visual feature extraction at scale, effectively supporting tasks like detection, segmentation and open-vocabulary understanding. Recent works have also introduced 2D foundation models into 3D by utilizing NeRF mildenhall2021nerf or 3DGS kerbl20233d. Existing 3DGS approaches fall into three categories: per-scene optimization-based zhou2024feature3DGS; zheng2024gaussiangrasper; Qin2024LangSplat; guo2024semantic; QU2024GOI; Zuo2024FMGS; peng2023openscene; Wu2024OpenGaussians, per-scene optimization-free joseph2024GWfeatureBP; marrie2024ludvig; cheng2024occam and generalizable li2025SceneSplat7k. Per-scene optimization-based LGS emerged first, with LERF kerr2023lerf; rashid2023language and LangSplat Qin2024LangSplat integrating 2D foundation model features into NeRF- and GS-based models, a design adopted by many later methods. Per-scene optimization-free LGS, such as OccamLGS cheng2024occam, LUDVIG marrie2024ludvig and Gradient Weighted 3DGS joseph2024GWfeatureBP further optimized this process by employing feature lifting, directly projecting all 2D features into 3DGS, reducing computation time from hours to seconds. SceneSplat li2025SceneSplat7k is the first generalizable LGS method, trained and evaluated directly on 3D data, achieving significant gains in open-vocabulary scene understanding with substantially faster inference on novel scenes. However, most of these methods have been evaluated in constrained 2D settings, with a limited number of scenes and semantic classes. We therefore introduce a comprehensive benchmark and use it to systematically evaluate recent language Gaussian Splatting methods.
@@
3D Generalizability. Large-scale training and generalization in 2D have become increasingly mature Radford2021CLIP; tschannen2025siglip; lyu2025realrag; lyu2024unibind; ravi2024sam; ravi2024sam2; zhong2025omnisam. In contrast, achieving semantic and language generalization in 3D remains challenging and relatively underexplored chen2024splatformer; PointMAE; zhao2021point; wu2022point; wu2024ptv3; lee2025mosaic3d; wu2025sonata. One major obstacle is the limited quantity and fragmented nature of 3D datasets, as no single dataset provides the needed combination of (i) large-scale coverage, (ii) multi-scale diversity (from object-centric to indoor and outdoor scenes), and (iii) rich multimodal information (geometry, appearance, and semantics). As a result, even state-of-the-art 3D models struggle to generalize broadly. Mosaic3D-5.6M compiles over 30K RGB-D scenes with 5.6 million mask-text pairs by leveraging 2D vision-language models, yielding an significant resource for open-vocabulary 3D segmentation lee2025mosaic3d. The SceneSplat-7K dataset supports open-vocabulary training but with limited scale li2025SceneSplat7k. These efforts represent early steps toward 3D foundation models, aiming to enable robust, optimization-free, and generalizable 3D reasoning across vision-language tasks. Therefore, we introduce SceneSplat-49K, a large and comprehensive dataset designed to unlock scaling laws in generalizable 3DGS tasks.

### v4

@@ Abstract
3D Gaussian Splatting (3DGS) serves as a highly performant and efficient encoding of scene geometry, appearance, and semantics. Moreover, grounding language in 3D scenes has proven to be an effective strategy for 3D scene understanding. Current Language Gaussian Splatting line of work falls into three main groups: (i) per-scene optimization-based, (ii) per-scene optimization-free, and (iii) generalizable approach. However, most of them are evaluated only on rendered 2D views of a handful of scenes and viewpoints close to the training views, limiting their ability and insight into holistic 3D understanding. To address this gap, we propose the first large-scale benchmark that systematically assesses these three groups of methods directly in 3D space, evaluating on 1060 scenes across three indoor datasets and one outdoor dataset. SceneSplat-Bench results demonstrate a clear advantage of the generalizable paradigm, particularly in relaxing the scene-specific limitation, enabling fast feed-forward inference on novel scenes, and achieving superior segmentation performance. We further introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising around 49K diverse indoor and outdoor scenes obtained from multiple sources, with which we demonstrate the generalizable approach could harness strong data priors. Our codes, benchmark, and datasets are available at https://scenesplatpp.gaussianworld.ai/.
@@ 1 Introduction
The field of 3D computer vision is highly focused on capturing the geometry Witkin (1981); Kazhdan et al. (2006); Hirschmuller (2008); Newcombe et al. (2011); Schönberger and Frahm (2016); Park et al. (2019); Shi et al. (2023); Ren et al. (2024) and the visual appearance Grimson (1981); Kutulakos and Seitz (1999); Furukawa and Ponce (2010); Mildenhall et al. (2020); Kerbl et al. (2023b) of a scene, as well as understanding its content Charles et al. (2017); Wu et al. (2024a); Peng et al. (2023); Zhu et al. (2023); Sarkar et al. (2025). Recently, 3D Gaussian Splatting (3DGS) Kerbl et al. (2023b) has emerged as the most desirable 3D representation due to its unique ability to jointly encode geometry, appearance, and scene understanding properties Ye et al. (2024); Shi et al. (2024); Jiang et al. (2024) in a compact form that can be effectively optimized from 2D posed images Ye et al. (2025). Furthermore, vision-language reasoning represents the most promising direction for 3D scene understanding Rozenberszki et al. (2022); Ding et al. (2023); Yang et al. (2024a); Ma et al. (2025); Jiang et al. (2024); Wu et al. (2023); Halacheva et al. (2025), as it bridges the visual and geometric attributes of a scene with the language we use to define, describe, and reason about concepts. Therefore, this work focuses exclusively on vision-language scene understanding with 3DGS.
@@
The most relevant methods for language Gaussian Splatting (LGS) can be categorized into three groups. The first two groups begin by extracting 2D features from all training images using vision-language foundation models (e.g., CLIP Radford et al. (2021a)). The first group then performs gradient-based per-scene optimization Qin et al. (2024); Zuo et al. (2024); Zhou et al. (2024); Wu et al. (2024b); Qu et al. (2024), assigning feature vectors to each 3D Gaussian primitive and optimizing them such that their renderings align with the corresponding 2D feature maps. The second group also operates on a per-scene basis but adopts an optimization-free approach Joseph et al. (2024); Marrie et al. (2024); Cheng et al. (2024). Instead of iterative refinement, these methods directly lift 2D features to 3D primitives via weighted feature aggregation schemes. Lastly, we consider the generalizable approach. To date, only a single method has been trained on 2K indoor scenes Li et al. (2025), learning to directly predict a vision-language feature vector for each 3D Gaussian primitive. Although 2D vision-language features are required during training, inference relies solely on 3D Gaussians, eliminating the need to prepare additional 2D feature maps. Furthermore, all features are produced in a single forward pass.
@@
While these methods represent important advances in integrating vision-language reasoning with 3DGS, their evaluation protocols still suffer from several critical limitations. First, most methods are evaluated on only a small number of selected scenes Qin et al. (2024); Zuo et al. (2024); Zhou et al. (2024); Wu et al. (2024b); Qu et al. (2024); Joseph et al. (2024); Cheng et al. (2024). As a result, such evaluations carry a high risk of producing scene-specific rather than generalizable conclusions. Moreover, the absence of a standardized benchmark further limits systematic and comparable evaluation. Second, most methods are evaluated close to the training views Qin et al. (2024); Marrie et al. (2024); Wu et al. (2024b). Consequently, there is no guarantee that the results generalize well to novel viewpoints. Finally, most evaluations are conducted in 2D rather than directly in 3D space Qin et al. (2024); Zuo et al. (2024); Zhou et al. (2024); Wu et al. (2024b); Qu et al. (2024); Joseph et al. (2024); Marrie et al. (2024); Cheng et al. (2024). However, 3D scene understanding fundamentally concerns performance in 3D space, which cannot be fully inferred from 2D projections. Therefore, we introduce SceneSplat-Bench, an evaluation benchmark for 3DGS vision-language scene understanding that includes 1060 scenes spanning 325 unique semantic classes. It evaluates performance in 3D for each segment of the scene, and we use it to assess representative methods from all three groups, as presented in Tab.˜1.
@@
Among the evaluated approaches, the generalizable 3D scene understanding paradigm demonstrates the strongest potential. It eliminates the need for substantial per-scene computation at inference and enables a single forward pass per 3D scene. In addition, this approach enables the application of 3D foundation models for scene understanding without the need for task- or scene-specific retraining. This concept aligns with established practices in 2D computer vision, where pretrained foundation models have been widely adopted He et al. (2016); Dosovitskiy et al. (2021); Oquab et al. (2023); Ravi et al. (2024a); Radford et al. (2021a); Liu et al. (2023). Foundation models tailored to 3DGS have already emerged across key areas, including reconstruction Charatan et al. (2024); Chen et al. (2024); Zhang et al. (2024); Ziwen et al. (2024), scene understanding Li et al. (2025), and generation Xiang et al. (2024); Roessle et al. (2024). Lastly, benchmark results demonstrate the state-of-the-art performance of the generalizable approach, highlighting its capacity to leverage data priors and extract meaningful vision-language features—effectively mitigating the noise inherent in weak supervision.
@@
To facilitate the development of generalizable 3DGS scene understanding, we introduce SceneSplat-49K, a carefully curated 3DGS dataset comprising diverse indoor and outdoor scenes collected from multiple sources. Existing large-scale 3DGS datasets primarily focus on single objects Ma et al. (2024); Liu et al. (2024b); wei et al. (2025), whereas scene-level datasets Li et al. (2025) remain limited in scale. To the best of our knowledge, SceneSplat-49K represents the most extensive open-source dataset of complex, high-quality 3DGS reconstructions at the full-scene level. The preparation of the dataset required approximately 891 GPU days and considerable human involvement. Furthermore, we demonstrate that training a generalizable 3DGS scene understanding method on a larger subset from SceneSplat-49K leads to improved performance, achieving state-of-the-art results on the benchmark.
@@
Methods Generalizable Optimization- free Runtime / Scene Benchmark Dataset Scene Number Class Number
@@
LangSplat Qin et al. (2024) 621 min LERF Kerr et al. (2023), 3DOVS Liu et al. (2024a) 9 88
FMGS Zuo et al. (2024) 76 min LERF Kerr et al. (2023), 3DOVS Liu et al. (2024a) 11 123
Feature3DGS Zhou et al. (2024) 235 min Replica Straub et al. (2019) 4 7
OpenGaussian Wu et al. (2024b) 297 min LERF Kerr et al. (2023), ScanNet Dai et al. (2017)(10scenes) 14 76
GOI Qu et al. (2024) 252min Mip-NeRF360 Barron et al. (2021), Replica Straub et al. (2019) 8 30
@@
Gradient-Weighted 3DGS  Joseph et al. (2024) 4 min LERF Kerr et al. (2023), 3DOVS Liu et al. (2024a) 9 88
LUDVIG Marrie et al. (2024) 5.6 min LERF Kerr et al. (2023), SPIn-NeRF Mirzaei et al. (2023), NVOS Ren et al. (2022) 21 80
OccamLGS Cheng et al. (2024) 3.2 min LERF Kerr et al. (2023), 3DOVS Liu et al. (2024a) 9 88
@@
SceneSplat Li et al. (2025) 0.24 min ScanNet Dai et al. (2017), ScanNet++ Yeshwanth et al. (2023),Matterport3D Chang et al. (2017) 732 321
SceneSplat-Bench ScanNet Dai et al. (2017), ScanNet++ Yeshwanth et al. (2023), Matterport3D Chang et al. (2017), HoliCity Zhou et al. (2021), 1060 325
@@ 2 Related Works
Gaussian Splatting Datasets. Most existing 3DGS datasets remain object-centric, synthetic, or limited in scale and scope. Moreover, recent works Charatan et al. (2024); Roessle et al. (2024); Ziwen et al. (2024); Chen et al. (2025) showcase 3DGS generalization, yet do not fully release their datasets, thus limiting reproducibility and scalability. Among publicly available datasets, ShapeSplat Ma et al. (2024) transforms 65K object-level CAD models into 3DGS. 3DAffordSplat wei et al. (2025) focuses on functionality, providing 23K object instances labeled with affordances across 21 categories. uCO3D Liu et al. (2024b) expands the object-centric landscape with over 1,000 annotated categories and 360 coverage. Furthermore, scene-level 3DGS remains underexplored, with SceneSplat Li et al. (2025) being the only available dataset providing 6.8K indoor scenes to support open-vocabulary 3D understanding, though its scope is limited to the indoor domain. In summary, existing datasets fall short of enabling generalizable, multi-modal learning across diverse scenes. To address this, we introduce SceneSplat-49K, a large-scale dataset of 49K 3DGS scenes spanning indoor and outdoor environments, enriched with vision-language embeddings to support scene-level understanding.
@@
Open-Vocabulary Scene Understanding. Recent advancements in open-vocabulary models have significantly expanded the capabilities of vision-language understanding. Foundation models Zhang et al. (2022); Oquab et al. (2023); Kirillov et al. (2023); Ravi et al. (2024b); Radford et al. (2021b); Zhai et al. (2023); Tschannen et al. (2025) enabled visual feature extraction at scale, effectively supporting tasks like detection, segmentation and open-vocabulary understanding. Recent works have also introduced 2D foundation models into 3D by utilizing NeRF Mildenhall et al. (2021) or 3DGS Kerbl et al. (2023a). Existing 3DGS approaches fall into three categories: per-scene optimization-based Zhou et al. (2024); Zheng et al. (2024); Qin et al. (2024); Guo et al. (2024); Qu et al. (2024); Zuo et al. (2024); Peng et al. (2023); Wu et al. (2024b), per-scene optimization-free Joseph et al. (2024); Marrie et al. (2024); Cheng et al. (2024) and generalizable Li et al. (2025). Per-scene optimization-based LGS emerged first, with LERF Kerr et al. (2023); Rashid et al. (2023) and LangSplat Qin et al. (2024) integrating 2D foundation model features into NeRF- and GS-based models, a design adopted by many later methods. Per-scene optimization-free LGS, such as OccamLGS Cheng et al. (2024), LUDVIG Marrie et al. (2024) and Gradient Weighted 3DGS Joseph et al. (2024) further optimized this process by employing feature lifting, directly projecting all 2D features into 3DGS, reducing computation time from hours to seconds. SceneSplat Li et al. (2025) is the first generalizable LGS method, trained and evaluated directly on 3D data, achieving significant gains in open-vocabulary scene understanding with substantially faster inference on novel scenes. However, most of these methods have been evaluated in constrained 2D settings, with a limited number of scenes and semantic classes. We therefore introduce a comprehensive benchmark and use it to systematically evaluate recent language Gaussian Splatting methods.
@@
3D Generalizability. Large-scale training and generalization in 2D have become increasingly mature Radford et al. (2021a); Tschannen et al. (2025); Lyu et al. (2025, 2024); Ravi et al. (2024b, a); Zhong et al. (2025). In contrast, achieving semantic and language generalization in 3D remains challenging and relatively underexplored Chen et al. (2025); Pang et al. (2022); Zhao et al. (2021); Wu et al. (2022, 2024a); Lee et al. (2025); Wu et al. (2025). One major obstacle is the limited quantity and fragmented nature of 3D datasets, as no single dataset provides the needed combination of (i) large-scale coverage, (ii) multi-scale diversity (from object-centric to indoor and outdoor scenes), and (iii) rich multimodal information (geometry, appearance, and semantics). As a result, even state-of-the-art 3D models struggle to generalize broadly. Mosaic3D-5.6M compiles over 30K RGB-D scenes with 5.6 million mask-text pairs by leveraging 2D vision-language models, yielding an significant resource for open-vocabulary 3D segmentation Lee et al. (2025). The SceneSplat-7K dataset supports open-vocabulary training but with limited scale Li et al. (2025). These efforts represent early steps toward 3D foundation models, aiming to enable robust, optimization-free, and generalizable 3D reasoning across vision-language tasks. Therefore, we introduce SceneSplat-49K, a large and comprehensive dataset designed to unlock scaling laws in generalizable 3DGS tasks.